OpenStudy (anonymous):

can someone help me with this equation? 8p+6=7p+14

OpenStudy (aroub):

8p+6-7p-14=0 p-8=0 p=8

OpenStudy (aroub):

did u understand how did i do it?

OpenStudy (anonymous):

can you show me step by step

OpenStudy (aroub):

yes ofcourse , u have to find y so the first step u add all them to one side (its easier,but some people don't so as u like) then u see which numbers have the same variable and add them or numbers( like what u have hear in this example 8p and -7p which becomes p, and -14+6=-8)then next step u already simplified (p-8=0) the last step u have to put it as it was p=8(and when u move a number from side to the other side the sign of the number changes )

OpenStudy (aroub):

i really hope that helped! :)

OpenStudy (anonymous):

can u help me with another one please

OpenStudy (aroub):

yes sure what is it?

OpenStudy (anonymous):

-6m-7=-7m+8 this is to make sure im on the right tract

OpenStudy (aroub):

can u do it and ill check it for u so i could know if u understand please

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-6m-7=-7m+8 +6 +6 m-7=1 +1 +1 m=-6

OpenStudy (aroub):

i got m=15

OpenStudy (anonymous):

how

OpenStudy (keenan):

4/3

OpenStudy (aroub):

no wait ill show u : -6m-7=-7m+8 -6m-7+7m-8 m-15=0 m=15

OpenStudy (aroub):

did u understand?

OpenStudy (anonymous):

no because u didnt show me what u added r subtracted. is it -6m-7 and then 7m-8 and then u added them together

OpenStudy (keenan):

add like terms and then divide

OpenStudy (aroub):

i putted then in one side which becomes -6m-7+7m-8=0

OpenStudy (anonymous):

okay im gonna try that to see if i get the same answer.

OpenStudy (aroub):

try to put another question and try to solve it :)

OpenStudy (anonymous):

10y=2y+9+7y

OpenStudy (aroub):

it needs alot ofpracticing :)

OpenStudy (keenan):

combine like terms and divide

OpenStudy (anonymous):

okay

OpenStudy (aroub):

try u solve it please :)

OpenStudy (keenan):

who me

OpenStudy (anonymous):

no me

OpenStudy (anonymous):

10y=2y+9+7y= 10y+2y+7y+9=28

OpenStudy (anonymous):

is that correct

OpenStudy (keenan):

checking now

OpenStudy (aroub):

go on

OpenStudy (aroub):

no it wrong

OpenStudy (keenan):

yeah i got 9/19

OpenStudy (aroub):

i told u when u want to move numbers from side to side u change the sign

OpenStudy (anonymous):

okay ill try again

OpenStudy (anonymous):

10y=2y+7y+9= 10y-2y-7y+9 is that right so far

OpenStudy (keenan):

add all the y's together

OpenStudy (anonymous):

okay 10y+2y+7y=19y

OpenStudy (keenan):

good

OpenStudy (anonymous):

2 1-9

OpenStudy (anonymous):

is it right

OpenStudy (aroub):

very good :)

OpenStudy (keenan):

nope then divide 9 into 19 and you should get 9/19

OpenStudy (anonymous):

can i do the same thing with the 1st one

OpenStudy (keenan):

yes

OpenStudy (anonymous):

keenan, 9/19=2 1/9

OpenStudy (keenan):

got to go bye

OpenStudy (anonymous):

aroub r u there

OpenStudy (aroub):

yes sorry

OpenStudy (aroub):

good improvement :) hahah :)

OpenStudy (anonymous):

im gonna try and solve the next one and u can check it ,and thank you

OpenStudy (anonymous):

-7a+2+8a=5-25

OpenStudy (aroub):

yes ofcourse i will do it too and see if ur answer is right or not :)

OpenStudy (anonymous):

-7a-8a+2=5+25 did i rewrite this one right

OpenStudy (aroub):

why -8a ?

OpenStudy (aroub):

u didnt remove it from its place

OpenStudy (aroub):

so u dont change the sign

OpenStudy (anonymous):

oh,okay ;so -7a+8a+2=3 and 5+25=30

OpenStudy (anonymous):

a=10

OpenStudy (anonymous):

3/3=30/3 a=10

OpenStudy (aroub):

no i got a=18

OpenStudy (aroub):

no no what r u doing :( from where did u get 3 ???

OpenStudy (anonymous):

because -7a+8a=1 and 1+2=3

OpenStudy (anonymous):

so sorry

OpenStudy (aroub):

no! try again please

OpenStudy (anonymous):

is it -6 2/3

OpenStudy (aroub):

wait please i want to check i think ur right :)

OpenStudy (anonymous):

please don't be madd @ me i just dont know how to do this stuff. i fell off again

OpenStudy (aroub):

no no wait i got now 20

OpenStudy (anonymous):

ill be right back,please dont leave

OpenStudy (aroub):

wait please i want to check it one more time

OpenStudy (anonymous):

okay -7a+8a+2=3 5-25=-20 3/3=1a -20/3=a=-6 2/3

OpenStudy (anonymous):

is it right

OpenStudy (aroub):

-7a+2+8a=5-25 a+2=-20 a=-2-20 a=-22

OpenStudy (anonymous):

where did the -2 come from

OpenStudy (aroub):

i took it to the other side

OpenStudy (anonymous):

okay i got it

OpenStudy (anonymous):

thank you for not being upset with me right now i just get confused

OpenStudy (anonymous):

2(y+4)=3(y-2)

OpenStudy (aroub):

di u really got it? im ready to help more

OpenStudy (anonymous):

okay im getting it. you can help more i wont mind

OpenStudy (anonymous):

i have already started on the other one

OpenStudy (aroub):

ok me too i dont mind :)

OpenStudy (anonymous):

ok does y = 14

OpenStudy (aroub):

yes its true i will give u a medal for that :)

OpenStudy (anonymous):

thank you 4 my medal

OpenStudy (anonymous):

9(k+1)-(8k+2)=1

OpenStudy (aroub):

hahha u 2 thank you for the medal :)

OpenStudy (aroub):

ok u solve me check :)

OpenStudy (anonymous):

does k=7

OpenStudy (anonymous):

aroub ru there

OpenStudy (aroub):

umm... no its k=-6

OpenStudy (anonymous):

how

OpenStudy (aroub):

ok 9(k+1)-(8k+2)=1 9k+9-8k-2=1 k+7=1 k=-6

OpenStudy (aroub):

hahaha ok :)

OpenStudy (aroub):

can u do it right every step here please

OpenStudy (anonymous):

i finally got -6